# Adapting Web Agents with Synthetic Supervision

### v1

@@ 3.3 Trajectory Refinement
Refining a task during execution inevitably produces trajectories whose segments were collected under earlier task variants and thus misaligned with the final refined task $\tau^{\star}$. In addition, even without task edits, the agent may wander through dead-ends, repeat no-op actions (e.g., consecutive Scroll), or backtrack, naturally resulting in noisy trajectories for supervision.
To convert such trajectory $h$ into more clean and suitable one that is aligned well with $\tau^{\star}$, SynthAgent introduces a post-hoc offline trajectory refinement step after collection, without re-interacting the environment or fabricating new steps. Thanks to the global view of the full trajectory and finally defined task, we can consider a small set of conservative edits:
  1. (1)
    Remove(i)(i): delete step $(o_{i},a_{i})$ if it is task-irrelevant or redundant (e.g., repeated Scroll with identical observation).
  2. (2)
    Reorder(i,j)(i,j): swap two locally commutable steps $(o_{i},a_{i})$ and $(o_{j},a_{j})$ when their targets are independent (e.g., open filter then set sort).
  3. (3)
@@
  4. (4)
[…]
These edits are proposed by prompting the LLM with $(\tau^{\star},h_{T})$ and then applied to refine the trajectory into clean one $h_{T}^{\star}$. The refined pair $(\tau^{\star},h_{T}^{\star})$ is expected to be less noisy and better consistent, thus more suitable for fine-tuning. For ablations, we also retain the original trajectories to quantify the effect of post-hoc editing.
@@ 4.1 Experimental Setup
We conduct experiments in WebArena (zhou2024webarenarealisticwebenvironment), a fully controllable suite of five websites: e-commerce (Shopping), content management (CMS), social forum (Reddit), developer platform (Gitlab), and map navigation (Maps). The environment provides simplified html code (accessibility tree) and visual screenshots as observations, and accepts common browser interactions listed in Table 5. We deliberately choose it as our sole testbed for three reasons. First, compared with simplified environments such as WebShop (yao2022webshop), each website in WebArena exposes deep task structure (multi pages, authentication, and full functionality), which better matches the realistic scenarios. Second, WebArena provides Docker deployments, ensuring the reproducibility, in contrast to online benchmarks such as Mind2Web (deng2023mind2webgeneralistagentweb) where pages may change over time. Finally, since web agents extensively interact with browsers during both exploration and evaluation, running in a local environment keeps experimental cost manageable and avoids the operational risks of live-web interaction.
@@
For both data synthesis and evaluation, we set a maximum step budget of $30$ per episode, following recommendations from (zhou2024webarenarealisticwebenvironment). All methods are synthesizing up to $500$ task-trajectory pairs per website for agent fine-tuning. During execution and training, they are using the same prompt template and action space. We also use the same GPT-4.1 for both synthesis and refinement for SynthAgent. We use a context window of $3$ most recent steps for efficiency. For agent fine-tuning, we mix synthesized data from five websites to train a single model with a learning rate of 1e-5 and batch size of $32$ for $3$ epochs. For evaluation, we restart the environment for every run, and report the success rate of completing the task from test set. More details are in Appendix A.1.
@@
Table 1: Performance comparison across different methods and websites. “SFT” means directly use the tasks from test set paired with trajectories collected by GPT-4.1 for agent fine-tuning. The best performance among synthetic methods is highlighted in bold. In most cases, ours SynthAgent outperforms other methods with synthetic data.
@@ 4.2 Main Results
The overall performance of different methods is shown in Table 1. From the results, we have the following observations: (1) Training on test set tasks (SFT) provides an informative upper bound for the following experiments using synthetic data. (2) Self-Instruct struggles without environment grounding, showing limited improvements over base models. This indicates the importance of interacting with the environment for creating environment specific tasks. (3) Explorer performs surprisingly poorly, even degrading base model performance. We find that this poor performance is due to: first, a substantial portion of its collected trajectories (about $68.3\%$) exceed the execution step budget without completing the task, which will be discussed later. This is because continuous task refinement strategy frequently changes task goals and adds new details during execution, causing the agent to deviate from the original objective. Second, the original Explorer paper states that their method operates on web pages without authentication requirements. This limitation suggests that Explorer may be unsuitable for complex, realistic websites that involve complex functionality. (4) Our SynthAgent substantially improves web agent adaptation with fully synthetic supervision, consistently outperforming all baselines. Compared to base models and OS-Genesis, we achieve average absolute gains of +10.2+10.2 and +5.1+5.1, respectively. SynthAgent also shows closer gap to the upper bound SFT, demonstrating its effectiveness in synthesizing high-quality data. The success of SynthAgent over baselines highlights the importance of dual refinement in both task and trajectory synthesis, which ultimately mitigates hallucinations and noise, offering better agent fine-tuning.
@@ 5 Analysis
In this section, we conduct further analyses to understand the effectiveness of SynthAgent.
@@ 5.1 Synthetic Data Quality
Self-Instruct 49.2 - 69 6.3 56.6
OS-Genesis 39.8 - 83 5.1 54.1
Explorer 69.8 8.6 54 20.5 48.1
SynthAgent 70.2 2.0 95 7.5 92.5
w/o TR 68.2 - 86 7.0 78.1
w/o JR 69.5 2.0 95 8.8 86.3
Table 2: Statistics of synthesized data. Quality metric is rated on a scale of 0-100 for every sample. Diversity metric is measured by 100 samples as a whole. ALL these metrics are rated by GPT-4.1. TR and JR stand for task and trajectory refinement, respectively.
@@
Synthesized task diversity is the essential prerequisite for effective agent adaptation. Self-Instruct produces points heavily clustered in a small region of the embedding space, which reflects: without environment grounding, the LLM “imagined” tasks results in repetitive and narrow task variations. In contrast, OS-Genesis achieves better diversity ($83$) through random environment exploration. Explorer exhibits the poorest diversity ($54$) with a highly clustered pattern, indicating that its initial coarse-grained task proposals from the homepage constrain the space of possible tasks. SynthAgent achieves the highest diversity score after task refinement, with points closely resembling the distribution of human written tasks. This suggests that our categorized exploration and task refinement strategies effectively covers different functions and interaction depths.
@@
Table 2 shows that our method achieves the highest trajectory quality ($92.5$), substantially outperforming other methods. From the ablation, we can find the improvement is enabled by our dual refinement. First, task refinement during trajectory collection ensures tasks are feasible and grounded in actual observations. Second, trajectory refinement with global context effectively mitigates noise from execution and task edits. To understand why Explorer underperforms despite similar refinement, we analyze trajectory completion rates in Table 3. Explorer’s continuous task refinement during execution changes goals and adds details at each step, causing agents to deviate from original objectives, resulting in 68.3% of trajectories exceeding the step budget without task completion. In contrast, SynthAgent has 73.8% success rate with task refinement and 92.5% after trajectory refinement, demonstrating the advantage of the proposed dual refinement design. In addition, compared with $0.22\mathdollar$ cost of Explorer, SynthAgent only costs only about $60\%$ of it due to more efficient exploration and refinement.
@@ 5.2 Ablation Study
+$S$ 14.55 10.53 11.54 16.07 12.50 13.27
+CE 20.00 14.04 15.38 16.07 12.50 15.93
   +TR 18.18 12.28 11.54 19.64 21.88 16.81
     +JR 20.00 21.05 15.38 19.64 28.12 20.80
Table 4: Ablation study of SynthAgent. $S$ and CE stand for naive synthetic supervision and categorized exploration, respectively. TR and JR stand for task and trajectory refinement, respectively. Indentation indicates the addition of a new component based on the up-level.
To quantify the contribution of each component in SynthAgent, we conduct a systematic ablation study in Table 4. Starting from the base model, naive synthetic supervision (+SS) boosts it to $13.27\%$, showing data synthesis by simply interacting with the environment can help adaptation. Adding categorized exploration (+CE) significantly enhances performance of Shopping and Reddit websites, which need categorization to cover diverse functions from repetitive elements. Interestingly, adding only task refinement (+TR) yields a marginal $0.88$ improvement, which confirms that changing tasks during execution can introduce noise to trajectories. Thanks to the final trajectory refinement, +JR achieves the strongest performance at $20.80\%$. This big gap highlights the trajectory refinement is necessary for the success of prior task refinement, without which the contribution would be masked by noisy trajectories.
@@ 4. Current Page (only current view):
A.2 Task Generation

### v2

@@ 3.3 Trajectory Refinement
Task refinement during execution enables feasible task completion but introduces a side effect: trajectory segments collected under earlier task variants become misaligned with the final task $\tau^{\star}$. Additionally, even without task edits, agents may wander through dead-ends or repeat no-op actions. This motivates trajectory refinement as a post-hoc step that leverages the global view of the complete trajectory and the final task to remove noise. Critically, task and trajectory refinement are synergistic: the former ensures task feasibility while introducing noise, which the latter subsequently mitigates.
@@ 4.1 Experimental Setup
We conduct experiments in two benchmarks: (1) WebArena (zhou2024webarenarealisticwebenvironment), a suite of five websites: e-commerce (Shopping), content management (CMS), social forum (Reddit), developer platform (Gitlab), and map navigation (Maps), which are stable and controllable. (2) Online-Mind2web (onlinemind2web), an online benchmark spanning 136 live websites, which are more realistic and diverse. The environments provide simplified html code (accessibility tree) and screenshots as observations, and accept common browser interactions listed in Table 6. Generally, WebArena is more challenging due to its tasks often requiring authentication, thus involving more complex operations.
@@
For both data synthesis and evaluation, we set a maximum step budget of $30$ per episode, following recommendations from (zhou2024webarenarealisticwebenvironment). All methods are synthesizing up to $500$ task-trajectory pairs per website for agent fine-tuning. During execution and training, they are using the same prompt template and action space. We also use the same GPT-4.1 for both synthesis and refinement for SynthAgent. We use a context window of $3$ most recent steps for efficiency. For agent fine-tuning, we mix synthesized data from five websites to train a single model with a learning rate of 1e-5 and batch size of $32$ for $3$ epochs. More details are in Appendix A.1.
@@
Table 1: Performance comparison on WebArena benchmark. “SFT” means directly use the tasks from test set paired with trajectories collected by GPT-4.1 for agent fine-tuning. The best performance among synthetic methods is highlighted in bold. In most cases, ours SynthAgent outperforms other methods with synthetic data.
Method Judge GPT-4.1 GPT-5.1 WebJudge Avg.
GPT-4.1 28.00 18.67 27.00 24.56
Qwen 16.90 6.49 19.48 14.29
Self-Instruct 17.67 13.00 19.67 16.78
OS-Genesis 19.53 11.00 19.33 16.62
Explorer 8.33 3.67 13.33 8.44
SynthAgent 31.67 15.67 23.33 23.56
Table 2: Task success rate on the Online-Mind2Web benchmark under three different judges (GPT-4.1, GPT-5.1, and WebJudge (onlinemind2web)). Our method SynthAgent consistently outperforms baselines across judges, and is competitive with the GPT-4.1.
@@ 4.2 Main Results
The overall performance of different methods on WebArena is shown in Table 1. From the results, we have the following observations: (1) Training on test set tasks (SFT) provides an informative upper bound for the following experiments using synthetic data. (2) Self-Instruct struggles without environment grounding, showing limited improvements over base models. This indicates the importance of interacting with the environment for creating environment specific tasks. (3) Explorer performs surprisingly poorly, even degrading base model performance. This is primarily because its continuous refinement strategy produces overly long trajectories, $68.3\%$ of which exceed the step budget without task completion. This is because continuous task refinement strategy frequently changes task goals and adds new details during execution, causing the agent to deviate from the original objective. Additionally, the original Explorer operates on pages without authentication, limiting its applicability to complex websites, while our re-implementation added the authentication support for fair comparison (details in Appendix A.4). (4) Our SynthAgent substantially improves web agent adaptation with fully synthetic supervision, consistently outperforming all baselines. Compared to base models and OS-Genesis, we achieve average absolute gains of +10.2+10.2 and +5.1+5.1, respectively. SynthAgent also shows closer gap to the upper bound SFT, demonstrating its effectiveness in synthesizing high-quality data.
Generalization.
Beyond the controlled offline WebArena environments with only five websites, we further test the generalization of SynthAgent on Online-Mind2Web, without any synthesis for targeted websites or re-training in Table 2. We again observe that SynthAgent outperforms other baselines across three different judges and is even competitive with GPT-4.1. These results on both offline (WebArena) and online (Online-Mind2Web) benchmarks suggest that our comprehensive exploration and dual refinement design enables the agent to generalize well on diverse websites rather than overfitting to several fixed environments.
@@ 5 Analysis
In this section, we analyze the quality of synthesized data, validate each component’s contribution, and examine data scaling of SynthAgent.
@@ 5.1 Synthetic Data Quality
Self-Instruct 59.3 (51.8) - 69 (69) 6.3 41.7 (45.7)
OS-Genesis 56.9 (50.2) - 65 (76) 5.1 52.0 (55.6)
Explorer 73.1 (72.0) 8.6 46 (56) 20.5 36.4 (43.5)
SynthAgent 72.6 (71.5) 2.0 82 (88) 7.5 82.6 (83.4)
w/o TR 68.5 (69.1) - 84 (85) 7.0 70.2 (71.9)
w/o JR 72.0 (70.4) 2.0 82 (88) 8.8 68.7 (75.1)
Table 3: Statistics of synthesized data rated by GPT-5.1 (average across GPT-4.1, GPT-5.1, Qwen3-30B (yang2025qwen3) in parentheses). Quality is rated on a scale of 0-100. Diversity is measured by 100 samples as a whole. TR and JR stand for task and trajectory refinement. More results are in Appendix A.5.
@@
Synthesized task diversity is the essential prerequisite for effective agent adaptation. Self-Instruct produces points heavily clustered in a small region of the embedding space, which indicates, without real environment grounding, LLM “imagined” tasks results in narrow task variations. In contrast, OS-Genesis achieves better diversity ($83$) through random environment exploration. Explorer exhibits the poorest diversity ($54$) with a highly clustered pattern, indicating that its initial coarse-grained task proposals from the homepage constrain the space of possible tasks. SynthAgent achieves the highest diversity score after task refinement, with points closely resembling the distribution of human written tasks. This suggests that our categorized exploration and task refinement strategies effectively covers different functions and interaction depths. Detailed statistics and case study of categorized exploration are provided in Appendix A.2 and B, respectively.
@@
Table 3 shows that our method achieves the highest trajectory quality ($82.6$, avg. $83.4$ across judges), substantially outperforming other methods. The improvement is enabled by our dual refinement: task refinement ensures tasks are feasible and grounded in observations, while trajectory refinement with global context mitigates noise from execution and task edits. Notably, our trajectory refinement applies edits conservatively where reorder accounts for only $4.1\%$ of edit operations (vs. $39.3\%$ for remove), gated by explicit independence checks. A head-to-head evaluation on 100 reordered trajectories shows that reordered versions achieve higher win rates ($42\%$ vs. $27\%$) and quality scores ($68.9$ vs. $62.1$), confirming that constrained reordering improves trajectory quality without introducing instability. Detailed results are in Appendix A.3.
To understand why Explorer underperforms despite similar refinement, we analyze trajectory completion rates in Table 4. Explorer’s continuous refinement changes task intentions at each step, causing $68.3\%$ of trajectories to exceed the step budget. In contrast, SynthAgent achieves $96.5\%$ completion rate after trajectory refinement with only $60\%$ of Explorer’s OpenAI API cost.
@@ 5.2 Ablation Study
SynthAgent 20.00 21.05 15.38 19.64 28.12 20.80
w/o CE 18.18 14.04 15.38 16.00 25.00 17.26
w/o TR 16.37 14.04 15.38 16.07 18.75 15.93
w/o JR 18.18 12.28 11.54 19.64 21.88 16.81
w/o TR&JR 20.00 14.04 15.38 16.07 12.50 15.93
Table 5: Ablation study of SynthAgent. CE, TR, and JR stand for categorized exploration, task refinement, and trajectory refinement, respectively.
To quantify the contribution of each component in SynthAgent, we conduct an ablation study in Table 5. Removing categorized exploration (w/o CE) causes a $3.54\%$ drop, with notable degradation on CMS ($7.01\%$), confirming that systematic coverage of functional groups is essential for task diversity. Removing task refinement (w/o TR) leads to a $4.87\%$ drop, the largest among all components. This is because without correcting hallucinations during execution, many synthesized tasks can be infeasible. Removing trajectory refinement (w/o JR) results in a $3.99\%$ drop, indicating that post-hoc noise removal is crucial for producing clean supervision signals. Interestingly, removing both TR and JR (w/o TR&JR) on Shopping improves while on Maps degrades. This confirms that TR and JR are synergistic: TR introduces some noise to the trajectory due to task edits during execution which JR subsequently removes, and their combination achieves the best overall performance.
@@ A.1 Environment
Online-Mind2Web (onlinemind2web) contains 300 realistic and diverse tasks spanning 136 live websites. Each task is requiring the agent to interact with real websites. The evaluation is performed with LLM-as-a-Judge through key point identification, key screenshot identification, and final outcome judgment.
@@ 4. Current Page (only current view):
A.2 Categorized Exploration
To understand the robustness and consistency of our categorized exploration, we compute summary statistics across all pages visited during task synthesis. For each environment, we measure the average number of elements per page, the split between static and interactive elements, the number of categories, and the average number of elements per category in Table 7.
Env Avg. Ele Avg. Static Avg. Interact. Avg. Cat. Avg. Ele/Cat.
Shopping 115.5 62.3 53.3 7.9 6.7
CMS 84.2 37.6 46.6 6.6 7.1
Reddit 78.6 39.7 38.9 5.3 7.3
Gitlab 111.2 48.9 62.3 5.8 10.7
Maps 53.9 24.0 29.9 4.2 7.1
Overall 88.7 42.5 46.2 6.0 7.8
Table 7: Statistics of categorized exploration across all pages visited during task synthesis. “Avg. Ele” denotes the average number of DOM elements per page, “Avg. Static” and “Avg. Interact.” are the averages for non-interactive and interactive elements, respectively, “Avg. Cat.” is the average number of functional categories identified by the LLM, and “Avg. Ele/Cat.” is the average number of elements per category.
The relatively stable category granularity across environments suggests that the proposed categorized exploration behaves consistently despite large differences in page structures.
A.3 Trajectory Refinement Statistics
To understand the behavior of our trajectory refinement, we analyze the distribution of edit operations and the effect of reordering in Table 8.
Operation Remove Reorder Drop Keep
Count 982.5 41.8 11.0 1463.9
Percentage 39.3% 1.7% 0.4% 58.6%
Table 8: Distribution of trajectory refinement operations. Reorder accounts for only 4.1% of edit operations (Remove + Reorder), indicating conservative application.
Metric Original Reordered
Win Rate (%) 27.0 42.0
Tie Rate (%) 31.0 31.0
Quality Score 62.1 68.9
Table 9: GPT-5.1 head-to-head evaluation of 100 reordered trajectories. Reordered trajectories achieve higher win rates and quality scores.
The reorder operation is applied only when an explicit independence check between steps is satisfied, ensuring that swapping does not affect the environment state. Despite its conservative usage (4.1% of edits), reordering provides measurable quality improvements when applied.
A.4 Baseline Implementation
All baselines are re-implemented using GPT-4.1 for fair comparison. For Self-Instruct (wang2023selfinstructaligninglanguagemodels), we follow the standard prompting approach to generate tasks from seed examples, as similar implementation by sun2025osgenesisautomatingguiagent. For OS-Genesis (sun2025osgenesisautomatingguiagent), we adapt the generation part of the released codebase to WebArena, using random exploration to collect interaction triplets and generate tasks from single-step observations.
For Explorer (pahuja2025explorerscalingexplorationdrivenweb), we believe our implementation is a faithful adaptation of the released codebase with minimal changes for WebArena compatibility (e.g., aligning browser operations). We additionally implement authentication handling, which the original Explorer does not support, to allow operation on login-gated pages. We have carefully investigated Explorer’s poor performance: as shown in Table 4, the dominant failure mode arises from its overly aggressive refinement strategy ($8.6$ refinements per trajectory on average) that produces long trajectories and frequently exhausts the step budget ($68.3\%$ exceed the budget). Through trajectory inspection, we confirmed this behavior matches the algorithmic design rather than an implementation error.
A.5 Trajectory Quality Evaluation
To ensure our quality assessments are not dependent on a single model, we evaluate synthesized data using three different judges: GPT-4.1, GPT-5.1, and Qwen3-30B (yang2025qwen3). Results with GPT-5.1, GPT-4.1, and Qwen3-30B are shown in Table 3, Table 10, and Table 11, respectively.
Method Task Trajectory
Quality # Refine Diversity #Steps Quality
Self-Instruct 49.2 - 69 6.3 56.6
OS-Genesis 39.8 - 83 5.1 54.1
Explorer 69.8 8.6 54 20.5 48.1
SynthAgent 70.2 2.0 95 7.5 92.5
w/o TR 68.2 - 86 7.0 78.1
w/o JR 69.5 2.0 95 8.8 86.3
Table 10: Trajectory quality rated by GPT-4.1.
Method Task Trajectory
Quality # Refine Diversity #Steps Quality
Self-Instruct 46.8 - 68 6.3 38.9
OS-Genesis 53.9 - 80 5.1 60.7
Explorer 73.0 8.6 68 20.5 45.9
SynthAgent 71.8 2.0 88 7.5 75.0
w/o TR 70.5 - 85 7.0 67.5
w/o JR 69.8 2.0 88 8.8 70.3
Table 11: Trajectory quality rated by Qwen3-30B.
Across all three different judges, SynthAgent consistently produces higher-quality task-trajectory data compared to baselines. While absolute scores vary by judge, the relative ranking of methods remains stable, indicating that our improvements are robust across different evaluation models.
A.6 Task Generation
@@ Appendix B Case Study
Categorized Category Sets.
To further illustrate the functional breadth captured by our categorized categorization, we list representative category sets for each website:
  • Shopping: Global Navigation, Search and Discovery, Shopping Options / Filtering, Sorting and Viewing Options, Product Browsing, Product Details and Purchase, Cart and Comparison, User Account Management, Product Reviews & Feedback, Subscription.
  • CMS: Global Navigation, User Account Management, Order Management, Product Management, Tab Navigation, Reporting & Analytics, Search & Filters, Table Controls & Pagination, Admin Controls, Help and Information.
  • Reddit: Navigation, User Account & Profile, Notifications & Submissions, Content Submission, Content Interaction (Posts & Voting), Comments / Threading / Visibility, Moderation & Toolbox, Forum Subscription & Management, Search & Discovery, Footer / Site Info.
  • Gitlab: Menu & Utility Actions, User/Account Actions, Repository High-Level Actions, Search & Filtering, Commit History Navigation, Alerts and Notifications, Code Commenting/Review, Project Management, Navigation & Site Structure, Issue Management.
  • Maps: Authentication/Login, Account Registration, Map Editing, Navigation & Page Switching, Sharing, Map Control, Community and Support, Map Data, Directions Input & Management, Sidebar/Popup Controls.
Across these environments, our method consistently identifies meaningful, task-relevant functional groups (e.g., navigation, search, account management, content interaction), which supports the claim that categorized exploration is both expressive and robust enough to guide task synthesis beyond simple random exploration.